# Normalized Remainders: Definition, Origins, Backgrounds, Recent Results, and Basic Properties

### v1

@@ 4.2. Generating functions of Howard’s generalization of Stirling numbers of the second kind
The $r$-associate Stirling numbers of the second kind $S_{r}(k,\ell)$ for $k\geq\ell\geq 0$ and $r\in\mathbb{N}_{0}$ are defined [27, p. 303, Eq. (1.2)] by
@@
see also [61, Section 1.12]. The equation (4.4) can be reformulated as
@@
for $\ell,r\in\mathbb{N}_{0}$. The equation (3.8) is a special case $r=0$ of (4.5). Consequently, the integer powers $(T_{r}[\operatorname{e}^{x}])^{\ell}$ for $\ell,r\in\mathbb{N}_{0}$ can be regarded as the generating functions of the Stirling numbers of the second kind $S(k,\ell)$ and the $r$-associate Stirling numbers of the second kind $S_{r}(k,\ell)$ for $k\geq\ell\in\mathbb{N}_{0}$ and $r\in\mathbb{N}_{0}$. See also [21, Remark 4], [22, Remark 4], and [71, Remark 2].
@@ 4.3. Generating functions of Howard’s generalization of Stirling numbers of the first kind
The Stirling numbers of the first kind $s(n,k)$ for $n,k\in\mathbb{N}_{0}$ can be analytically generated [38, Theorem 3.14] by (3.17). The generating function $\bigl[\frac{\ln(1+x)}{x}\bigr]^{k}$ can be regarded as the power of the normalized remainder $(T_{0}[\ln(1+x)])^{k}$.
@@
in (4.7) is just the normalized remainder $T_{r}\bigr[\frac{\ln(1+x)}{x}\bigr]$ for $r\in\mathbb{N}_{0}$ or the normalized remainder $(T_{r}[\ln(1+x)])^{\ell}$. See also [21, Remarks 5 and 8] and [22, Remarks 5 and 8].
@@ 4.5. Generating function of Broder–Carlitz’s generalization of Stirling numbers of the second kind
Theorem 16 in [9] and Eq. (3.9) in [11] states that the $r$-Stirling numbers of the second kind for $r\in\mathbb{N}_{0}$ have the exponential generating function
@@ 5. Basic properties of normalized remainders
In this section, we recite the following basic properties of normalized remainders.
Theorem 5.1 ([50, Theorem 1]).
@@
Theorem 5.3 ([50, Theorem 1]).
If $f^{(i)}(x_{0})=0$ for $0\leq i<k$, but $f^{(k+1)}(x_{0})\neq 0$, and the normalized remainder $T_{n}[f(x),x_{0}]$ exists for some $n\geq k$, then the normalized remainder $T_{n}[f(x),x_{0}]$ satisfies
@@
6. Problems and guesses on normalized remainders
@@ 6.1. Problems on normalized remainders
The following are six problems on concrete normalized remainders.
@@ 6.1.5. Fifth problem
The integral representation (3.13) and the ideas with techniques in [52, Section 7] should be useful when studying this problem.
@@ 6.1.6. Sixth problem
6.2. Guesses on normalized remainders
The following are nine guesses on concrete normalized remainders.
6.2.1. First guess
In [36, Remark 4.1], we proposed the guess: The ratio $\frac{T_{2n+1}[\tan^{2}x]}{T_{2n-1}[\tan^{2}x]}$ for $n\in\mathbb{N}$ is concave in $x\in\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
6.2.2. Second guess
In the paper [65, Section 4.1], we posed the following guesses: For $n\in\mathbb{N}$, the ratio $\frac{T_{2n-1}[\tan^{2}x]}{T_{2n+1}[\tan^{2}x]}$ is completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$ and is absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
6.2.3. Third guess
In the paper [65, Section 4.2], we posed the following guesses: For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is a logarithmically absolutely monotonic function in $x\in\mathbb{R}$.
6.2.4. Fourth guess
In the paper [65, Section 4.2], we posed the following guesses: For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{n}[\operatorname{e}^{x}]}{T_{n+1}[\operatorname{e}^{x}]}$ is an absolutely monotonic function in $x\in\mathbb{R}$.
6.2.5. Fifth guess
In [47, Remark 7], we guessed that the functions $G_{n}(x)$ and $H_{n}(x)$ for $n\in\mathbb{N}$ are absolutely monotonic in $x\in\mathbb{R}$, where
$G_{n}(x)=\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n+1}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}-\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n+1}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n}{n}}\frac{x^{j}}{j!}$
@@ 6.2.5. Fifth guess
$H_{n}(x)=\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n}{n}}\frac{x^{j}}{j!}-\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}\Biggr]^{2}.$
This guess is stronger than the inequalities (3.11) and (3.12).
6.2.6. Sixth guess
Since the quantity $\mathcal{Q}(q,2\ell)$ for $q,\ell\in\mathbb{N}$ is positive, we guessed in [50, Section 1] that the normalized remainder $T_{2n+1}\bigl[\bigl(\frac{\arcsin x}{x}\bigr)^{q}\bigr]$ for $q\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$ is logarithmically absolutely monotonic in $x\in[0,1)$.
6.2.7. Seventh guess
6.2.8. Eighth guess
6.2.9. Ninth guess
@@ 7. Conclusions
Normalized remainders represent a novel concept in mathematical analysis and combinatorial number theory. This concept is grounded in well-established mathematical knowledge and supported by a rich historical foundation. The main results reviewed in this work serve as initial steps and a fundamental basis for further exploration. These findings illustrate that the study of normalized remainders is both intriguing and significant, with promising prospects for future development.

### v2

@@ 4.2. Generating functions of Howard’s generalization of Stirling numbers of the second kind
The $r$-associate Stirling numbers of the second kind $S_{r}(k,\ell)$ for $k\geq\ell\geq 0$ and $r\in\mathbb{N}_{0}$ are defined [28, p. 303, Eq. (1.2)] by
@@
see also [64, Section 1.12]. The equation (4.4) can be reformulated as
@@
for $\ell,r\in\mathbb{N}_{0}$. The equation (3.8) is a special case $r=0$ of (4.5). Consequently, the integer powers $(T_{r}[\operatorname{e}^{x}])^{\ell}$ for $\ell,r\in\mathbb{N}_{0}$ can be regarded as the generating functions of the Stirling numbers of the second kind $S(k,\ell)$ and the $r$-associate Stirling numbers of the second kind $S_{r}(k,\ell)$ for $k\geq\ell\in\mathbb{N}_{0}$ and $r\in\mathbb{N}_{0}$; see also [21, Remark 4], [22, Remark 4], and [74, Remark 2].
@@ 4.3. Generating functions of Howard’s generalization of Stirling numbers of the first kind
The Stirling numbers of the first kind $s(n,k)$ for $n,k\in\mathbb{N}_{0}$ can be analytically generated [39, Theorem 3.14] by (3.19). The generating function $\bigl[\frac{\ln(1+x)}{x}\bigr]^{k}$ can be regarded as the power of the normalized remainder $(T_{0}[\ln(1+x)])^{k}$.
@@
in (4.7) is just the normalized remainder $T_{r}\bigr[\frac{\ln(1+x)}{x}\bigr]$ for $r\in\mathbb{N}_{0}$ or the normalized remainder $(T_{r}[\ln(1+x)])^{\ell}$; see also [21, Remarks 5 and 8] and [22, Remarks 5 and 8].
@@ 4.5. Generating function of Broder–Carlitz’s generalization of Stirling numbers of the second kind
Theorem 16 in [9] and Eq. (3.9) in [11] state that the $r$-Stirling numbers of the second kind for $r\in\mathbb{N}_{0}$ have the exponential generating function
@@
Remark 4.1.
Some contents of Sections 4.1 to 4.5 were also posted at the websites https://mathoverflow.net/q/508432 (accessed on 23 February 2026) and https://math.stackexchange.com/q/5126180 (accessed on 25 February 2026).
@@ 5. Basic properties of normalized remainders
In this section, we recite some basic properties of normalized remainders.
Theorem 5.1 ([53, Theorem 1]).
@@
Theorem 5.3 ([53, Theorem 1]).
If $f^{(i)}(x_{0})=0$ for $0\leq i<k$ and the normalized remainder $T_{n}[f(x),x_{0}]$ exists for some $n\geq k$, but $f^{(k+1)}(x_{0})\neq 0$, then the normalized remainder $T_{n}[f(x),x_{0}]$ satisfies
@@
Theorem 5.4.
If $f(x)$ is an absolutely monotonic function on an interval $0,a]\subset[0,\infty)$ and its normalized remainder $T_{n}[f(x)]$ exists for some $n\in\mathbb{N}_{0}$, then its normalized remainder $T_{n}[f(x)]$ for some $n\in\mathbb{N}_{0}$ is also an absolutely monotonic function on $0,a]\subset[0,\infty)$.
Proof.
This follows from combining Definition 1.1 with a statement in [76, p. 156] which reads that, if $f(x)$ is absolutely monotonic in $[0,a]$ and if $f(x)=\sum_{n=0}^{\infty}A_{n}x^{n}$ for $x\in[0,a]$, then all coefficients $A_{n}$ are non-negative. ∎
6. Guesses and problems on normalized remainders
@@
6.1. Guesses on normalized remainders
The following are nine guesses on concrete normalized remainders.
6.1.1. First guess
In [37, Remark 4.1], we proposed the guess: The ratio $\frac{T_{2n+1}[\tan^{2}x]}{T_{2n-1}[\tan^{2}x]}$ for $n\in\mathbb{N}$ is concave in $x\in\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
6.1.2. Second guess
In the paper [68, Section 4.1], we posed the following guesses: For $n\in\mathbb{N}$, the ratio $\frac{T_{2n-1}[\tan^{2}x]}{T_{2n+1}[\tan^{2}x]}$ is completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$ and absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
6.1.3. Third guess
In the paper [68, Section 4.2], we posed the following guesses: For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is a logarithmically absolutely monotonic function in $x\in\mathbb{R}$.
6.1.4. Fourth guess
In the paper [68, Section 4.2], we posed the following guesses: For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{n}[\operatorname{e}^{x}]}{T_{n+1}[\operatorname{e}^{x}]}$ is an absolutely monotonic function in $x\in\mathbb{R}$.
6.1.5. Fifth guess
In [49, Remark 7], we guessed that the functions
$G_{n}(x)=\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n+1}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}-\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n+1}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n}{n}}\frac{x^{j}}{j!}$ (6.1)
and
$H_{n}(x)=\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n}{n}}\frac{x^{j}}{j!}-\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}\Biggr]^{2}$ (6.2)
for $n\in\mathbb{N}$ are absolutely monotonic in $x\in\mathbb{R}$.
The absolute monotonicity of $G_{n}(x)$ and $H_{n}(x)$ are stronger than the inequalities (3.11) and (3.12).
In [23, Section 1], we guessed that the function
$J_{n}(x)=\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n+1}}\frac{x^{j}}{j!}\Biggr]^{2}-\frac{n+1}{n+2}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+3}{n+2}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}$ (6.3)
for $n\in\mathbb{N}_{0}$ is absolutely monotonic in $x\in\mathbb{R}$,
6.1.6. Sixth guess
Since the quantity $\mathcal{Q}(q,2\ell)$ for $q,\ell\in\mathbb{N}$ is positive, we guessed in [53, Section 1] that the normalized remainder $T_{2n+1}\bigl[\bigl(\frac{\arcsin x}{x}\bigr)^{q}\bigr]$ for $q\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$ is logarithmically absolutely monotonic in $x\in[0,1)$.
6.1.7. Seventh guess
6.1.8. Eighth guess
6.1.9. Ninth guess
Remark 6.1.
The first five guesses mentioned above were also posted at the websites https://mathoverflow.net/q/508432 (accessed on 23 February 2026) and https://math.stackexchange.com/q/5126180 (accessed on 25 February 2026).
@@ 6.2. Problems on normalized remainders
The following are eight problems related to normalized remainders.
@@ 6.2.5. Fifth problem
The integral representation (3.13) and the ideas with techniques in [55, Section 7] should be useful when studying this problem.
@@ 6.2.6. Sixth problem
6.2.7. Seventh problem
It is well-known that, if a function $f(x)$ is of the period $2\ell>0$ on $\mathbb{R}$ and the integrals
$a_{n}=\frac{1}{\ell}\int_{-\ell}^{\ell}f(x)\cos\frac{n\pi x}{\ell}\operatorname{d\!}x,\quad n\in\mathbb{N}_{0}$
@@ 6.2.7. Seventh problem
$b_{n}=\frac{1}{\ell}\int_{-\ell}^{\ell}f(x)\sin\frac{n\pi x}{\ell}\operatorname{d\!}x,\quad n\in\mathbb{N}$
exist, then the trigonometric series
$\frac{a_{0}}{2}+\sum_{n=1}^{\infty}\biggl(a_{n}\cos\frac{n\pi x}{\ell}+b_{n}\sin\frac{n\pi x}{\ell}\biggr)$
is called the Fourier series of the function $f(x)$.
Similar to Definition 1.1, consider the function
$F_{n}[\ell,f(x)]=\frac{\frac{a_{0}}{2}+\sum_{k=1}^{n}\bigl(a_{k}\cos\frac{k\pi x}{\ell}+b_{k}\sin\frac{k\pi x}{\ell}\bigr)}{a_{n+1}\cos\frac{(n+1)\pi x}{\ell}+b_{n+1}\sin\frac{(n+1)\pi x}{\ell}}$
for $n\in\mathbb{N}$ and $x\in\mathbb{R}$ and discuss its properties.
6.2.8. Eighth problem
Let $\boldsymbol{s}_{n}=(s_{1},s_{2},\dotsc,s_{n})$ and $\boldsymbol{t}_{n}=(t_{1},t_{2},\dotsc,t_{n})\in\mathbb{R}^{n}$. A $n$-tuple $\boldsymbol{s}_{n}$ is said to strictly majorize $\boldsymbol{t}_{n}$, denoted by $\boldsymbol{t}_{n}\prec\boldsymbol{s}_{n}$ or $\boldsymbol{s}_{n}\succ\boldsymbol{t}_{n}$, if
$\bigl(s_{[1]},s_{[2]},\dotsc,s_{[n]}\bigr)\neq\bigl(t_{[1]},t_{[2]},\dotsc,t_{[n]}\bigr),\qquad\sum_{i=1}^{n}s_{i}=\sum_{i=1}^{n}t_{i},$
and
$\sum_{i=1}^{k}s_{[i]}>\sum_{i=1}^{k}t_{[i]},\quad 1\leq k\leq n-1$
are valid, where
$s_{[1]}\geq s_{[2]}\geq\dotsm\geq s_{[n]}\quad\text{and}\quad t_{[1]}\geq t_{[2]}\geq\dotsm\geq t_{[n]}$
are rearrangements of $\boldsymbol{s}_{n}$ and $\boldsymbol{t}_{n}$ in a descending order; see [40, p. 8, Defnition A.1].
In [23, Section 5], the function $G_{n}(x)$ defined by (6.1) was generalized as
$G_{m,n;k,\ell}(x)=-\begin{vmatrix}S_{m,n}(x)&S_{m,n+\ell}(x)\\ S_{m+k,n}(x)&S_{m+k,n+\ell}(x)\end{vmatrix}$ (6.4)
for $m,n\in\mathbb{N}_{0}$ and $k,\ell\in\mathbb{N}$, where $S_{m,n}(x)$ is defined by (3.14).
In [23, Section 5], basing on an inequality in [23, Theorem 7], we defined
$K_{n}(x)=\frac{n+1}{n+2}S_{1,n+2}(x)S_{1,n}(x)-\frac{n}{n+1}S_{1,n+1}^{2}(x)$ (6.5)
for $n\in\mathbb{N}_{0}$ and $x\in\mathbb{R}$.
The functions $H_{n}(x)$ defined by (6.2), $J_{n}(x)$ defined by (6.3), and $K_{n}(x)$ defined by (6.5) can be generalized as follows:
  1. 1)
    For $n\in\mathbb{N}_{0}$ and $(s_{1},s_{2}),(t_{1},t_{2})\in\mathbb{N}_{0}^{2}$ such that $(s_{1},s_{2})\succ(t_{1},t_{2})$, define
    $H_{n;\boldsymbol{s}_{2},\boldsymbol{t}_{2}}(x)=S_{s_{1},n+1}(x)S_{s_{2},n+1}(x)-S_{t_{1},n+1}(x)S_{t_{2},n+1}(x).$ (6.6)
  2. 2)
    For $m\in\mathbb{N}_{0}$ and $(s_{1},s_{2}),(t_{1},t_{2})\in\mathbb{N}_{0}^{2}$ such that $(s_{1},s_{2})\succ(t_{1},t_{2})$, define
    $J_{m;\boldsymbol{s}_{2},\boldsymbol{t}_{2}}(x)=S_{m,t_{1}}(x)S_{m,t_{2}}(x)-A(m;s_{1},s_{2};t_{1},t_{2})S_{m,s_{1}}(x)S_{m,s_{2}}(x).$ (6.7)
  3. 3)
    For $m\in\mathbb{N}_{0}$ and $(s_{1},s_{2}),(t_{1},t_{2})\in\mathbb{N}_{0}^{2}$ such that $(s_{1},s_{2})\succ(t_{1},t_{2})$, define
    $K_{m;\boldsymbol{s}_{2},\boldsymbol{t}_{2}}(x)=S_{m,s_{1}}(x)S_{m,s_{2}}(x)-B(m;s_{1},s_{2};t_{1},t_{2})S_{m,t_{1}}(x)S_{m,t_{2}}(x).$ (6.8)
Problem 3 in [23] stated that,
  1. 1)
    for $m,n\in\mathbb{N}_{0}$ and $k,\ell\in\mathbb{N}$, the function $G_{m,n;k,\ell}(x)$ defined by (6.4) is absolutely monotonic in $x\in\mathbb{R}$;
  2. 2)
    for $n\in\mathbb{N}_{0}$ and $(s_{1},s_{2}),(t_{1},t_{2})\in\mathbb{N}_{0}^{2}$ such that $(s_{1},s_{2})\succ(t_{1},t_{2})$, the function $H_{n;\boldsymbol{s}_{2},\boldsymbol{t}_{2}}(x)$ defined by (6.6) is absolutely monotonic in $x\in\mathbb{R}$;
  3. 3)
    for $m\in\mathbb{N}_{0}$ and $(s_{1},s_{2}),(t_{1},t_{2})\in\mathbb{N}_{0}^{2}$ such that $(s_{1},s_{2})\succ(t_{1},t_{2})$, find the best constants $0<A(m;s_{1},s_{2};t_{1},t_{2})<1$ and $0<B(m;s_{1},s_{2};t_{1},t_{2})<1$ such that the functions $J_{m;\boldsymbol{s}_{2},\boldsymbol{t}_{2}}(x)$ defined by (6.7) and $K_{m;\boldsymbol{s}_{2},\boldsymbol{t}_{2}}(x)$ defined by (6.8) are absolutely monotonic in $x\in\mathbb{R}$.
@@ 7. Conclusions
Normalized remainders represent a novel concept in mathematical analysis and combinatorial number theory. The concept is grounded in well-established mathematical knowledge and supported by a rich historical foundation. The main results reviewed in this work serve as initial steps and a fundamental basis for further exploration. These findings illustrate that the study of the normalized remainders is both intriguing and significant, with promising prospects for future development.
Remark 7.1.
This chapter is a slightly extended and modified version of the electronic arXiv preprint [51].
Acknowledgements
The author is grateful to anonymous referees for their careful comments and helpful suggestions to the original version of this chapter.